A table of values of a linear function is shown below. Find the output when the input is n. Input: 1,2,3,4,n Output:7,11,15,19

Answer by greenestamps(13200)



We are looking for an equation of the form

y = ax+b

Each time the input increases by 1, the output increases by 4. That means a is 4. So our equation is

y = 4x+b

When the input x is 1, the output y is 7:

7 = 4(1)+b
7 = 4+b
b = 3

The linear equation is y = 4x+3.